# Evaluating the Elementary Multilingual Capabilities of Large Language Models with MultiQ

### v1

@@ Abstract
Large language models (LLMs) need to serve everyone, including a global majority of non-English speakers. However, most LLMs today, and open LLMs in particular, are often intended for use in just English (e.g. Llama2, Mistral) or a small handful of high-resource languages (e.g. Mixtral, Qwen). Recent research shows that, despite limits in their intended use, people prompt LLMs in many different languages. Therefore, in this paper, we investigate the basic multilingual capabilities of state-of-the-art open LLMs beyond their intended use. For this purpose, we introduce MultiQ, a new silver standard benchmark for basic open-ended question answering with 27.4k test questions across a typologically diverse set of 137 languages. With MultiQ, we evaluate language fidelity, i.e. whether models respond in the prompted language, and question answering accuracy. All LLMs we test respond faithfully and/or accurately for at least some languages beyond their intended use. Most models are more accurate when they respond faithfully. However, differences across models are large, and there is a long tail of languages where models are neither accurate nor faithful. We explore differences in tokenization as a potential explanation for our findings, identifying possible correlations that warrant further investigation.
@@ 1 Introduction
Languages other than English remain underrepresented and underserved by state-of-the-art language technologies, posing a barrier to equal and inclusive AI (Bender, 2011; Joshi et al., 2020). While proprietary large language models (LLMs) like GPT-4 (OpenAI, 2023) may answer questions and follow instructions in many different languages, even the best and most popular open LLMs are much more restricted in their language coverage: Llama2-chat (Touvron et al., 2023b), for example, is “intended for commercial and research use in English”. [^1] https://huggingface.co/meta-llama/Llama-2-7b-hf Yi is “bilingual” in English and Chinese [^2] https://huggingface.co/01-ai/Yi-34B-Chat, and Mistral-7b-instruct (Jiang et al., 2023) “only works in English”. [^3] https://mistral.ai/news/la-plateforme/
@@
Even though most open LLMs are restricted in their intended use to one or a handful of languages, datasets of real-world LLM usage show that people prompt LLMs in many different languages, often beyond their intended use (Zheng et al., 2023; Ouyang et al., 2023; Zhao et al., 2024). This has motivated initial research into the multilingual capabilities of monolingual models (Armengol-Estapé et al., 2022; Lai et al., 2023a). However, this research has mostly focused on older proprietary LLMs and on a relatively small number of languages and/or specific tasks.
In this paper, we investigate the basic multilingual capabilities of a variety of state-of-the-art chat-optimized open LLMs across a typologically diverse set of 137 languages. Specifically, we ask two main research questions that correspond to two dimensions of capability: 1) What is the multilingual language fidelity of current chat-optimized open LLMs, and 2) What is the multilingual question answering (QA) accuracy of current chat-optimised open LLMs? Language fidelity describes the ability to respond to prompts in the prompted language. QA accuracy describes the ability to give correct answers to open-ended questions, in the prompted language or any other. An ideal multilingual model would give answers that are both faithful and correct.
@@
    Interaction: Increased language fidelity may positively impact answer accuracy, given that model answers in the same language as prompted tend to be more accurate.
@@
    Explanation: With regard to the tokenization strategy of models for different languages, models tend to achieve higher accuracy on languages they can tokenize into subwords instead of characters or ASCII tokens.
@@ 2.1 Dataset Creation
For the initial English questions, we used two different sources to increase question diversity. a) We collected 100 questions from the LMSYS-Chat-1M dataset (Zheng et al., 2023), which catalogs real-world user interactions with LLMs. Specifically, we sampled all single-sentence English-language and sorted them by frequency. Then, we manually selected from the top until we reached 100 questions. This portion of our data directly reflects real-world LLM usage. b) We manually created another set of 100 questions evenly spread across 10 different subjects at elementary to middle school level (e.g.   mathematics and geography). To maximize the diversity of the questions, we prompted GPT-4 to provide us with a set of simple and clear questions with simple and clear answers for each of the subjects. We then iterated and manually selected questions until we reached 10 questions per subject. This portion of our data expands MultiQ’s topical coverage.
@@
Source Example Prompt
@@
LMSYS Was the year 2000 a leap year?
What is 2 + 2 * 3?
How many feet does a chicken have?
Table 1: Examples of English questions covered in MultiQ. We present three prompts from each source (GPT-4 and LMSYS) covering different domains.
@@ 2.2 Validation of Translations
We asked native speakers to annotate the correctness of the 200 translated MultiQ questions for 10 languages: Chinese, French, Farsi, German, Indonesian, Italian, Japanese, Russian, Spanish, and Urdu. Across these languages, annotators marked an average of 94.4% of translations as correct. Translations were least accurate for Indonesian and Farsi, at 89.0%, while the translations for Italian and German were the most accurate, at 99.0%. We present the full results in the Appendix A. Qualitatively, several annotators stated that some translations, while accurate in content, tended to be literal, word-for-word translations rather than natural expressions. Overall, the automated translation introduces some noise into MultiQ, but our validation results suggest that the amount of noise is limited. This is why we frame MultiQ as a silver standard benchmark that can provide meaningful insights into basic multilingual capabilities, even if exact results on individual test cases may not be perfectly reliable.
@@ 2.3 Typological Diversity
To estimate language distances, we use the lang2vec toolkit (Littell et al., 2017) which contains precomputed language distances based on the typological language information queried from the URIEL knowledge base  [^5] https://www.cs.cmu.edu/~dmortens/projects/7_project/. Following Ploeger et al. (2024), we calculate the distances between all pairs of languages in our dataset that have at least 5% coverage in the URIEL vectors. Figure 2 shows the distribution of pairwise geographic, syntactic, and genetic distances of all covered languages. We find that the languages in MultiQ cover a wide range of typologically similar and distant language pairs with an expected high skewness of genetic distance complementing previous research (Ploeger et al., 2024). We can therefore confidently speak of a high typological diversity of our dataset.
@@
Next, we calculate the typological feature coverage of the 137 languages in MultiQ, using language features provided in the Grambank database (Skirgård et al., 2023). For this purpose, we map the language IDs in MultiQ (obtained from the Google Translate API) to the Glottoids used by Grambank. [^6] For 14 languages in our dataset there is no matching entry in Grambank. To avoid incorrect mapping, we do not assign them manually and exclude these languages from the calculation Taken together, the languages in MultiQ cover at least 95.38% of the typological features recoded in Grambank. This underlines the typological diversity of our dataset.
@@ 3 Experiments and Results
Using MultiQ, we can now answer our two main research questions regarding the multilingual language fidelity and multilingual QA accuracy of current chat-optimized open LLMs. We first separately assess fidelity and accuracy, and then evaluate their relationship. [^9] We publish all data and code at https://github.com/paul-rottger/multiq
@@ 3.1 Overall Experimental Setup
We test six open-access LLMs that are both popular and competitive in performance with other state-of-the-art models as measured on standard (English-language) benchmarks such as the LMSys Leaderboard  [^10] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard and AlpacaEval. [^11] https://tatsu-lab.github.io/alpaca_eval/ Specifically, the 7B, 13B and 70B versions of Llama2-Chat (Touvron et al., 2023b), the 7B Mistral-Instruct-v0.1 (Jiang et al., 2023), the 8x7B Mixtral-Instruct-v0.1 (Jiang et al., 2024) and the 7B Qwen1.5-Chat model (Bai et al., 2023). We test three sizes of Llama2 to evaluate scaling on MultiQ. Llama2 and Mistral are explicitly intended for English use only, whereas Mixtral and Qwen are explicitly multilingual: Mixtral “handles English, French, Italian, German and Spanish”, while Qwen offers unspecified “multilingual support”.
@@ 3.2 Language Fidelity
Next, we conduct a more fine-grained analysis of model answers in “another language”. For this, we focus on the three models responding in another language more than 10% of the time (Qwen, Mistral, and Mixtral). Of these, Mistral demonstrates the highest level of diversity among the languages it responds in. As opposed to Mixtral and Qwen, which draw from a pool of around 240 languages, GlotLID identifies more than 360 distinct languages in Mistral’s responses. We take a look into the correspondence between the language family of the input and output language. While Mixtral provides a response in the same language family at least 68.7% of the time, if not in an exact match or responding in English, Qwen demonstrates this behavior only 60.6% of the time, while Mistral does so for 54.6%. Note, however, that for 6% of the answers of Mistral and 7% of the answers of Mixtral, the language family of the response languages could not be determined and is thus classified as ’unknown’. Additionally, the distribution of language families in MultiQ is highly skewed, with 45% of the languages belonging to the ’Indo-European’ family, which spans a very broad range of languages from Irish to Turkish.
@@
Finally, we analyze the relationship between the language of the prompt and the frequency of the respective response language of the models. Figure 4 shows correlation matrices for Mistral and Qwen. Corresponding matrices for the other models are shown in Appendix C. We find that Hindi is the most frequently selected language outside of the input language or English, closely followed by Indonesian. For Qwen, for example, these languages make up 20.4% of the “another language” category. This is visible by the thin yellow vertical lines leading to the axis ticks ind and hin in the Figure for both models respectively. It appears that the models lack support for numerous languages from India and Indonesia, thus treating Hindi and Indonesian as some kind of ’fallback’ languages for the wider language area. For example, the models often respond to languages such as Malay and Javanese in Indonesian, and languages such as Maithili, Konkani, and Bhojpuri in Hindi. However, this phenomenon can also be seen in smaller European languages. For example, the three models respond to questions in Croatian mostly in Italian, to those in Luxembourgish mostly in German, and those in Galician mostly in Portuguese. These observations underline the importance of improving multilingual models in the language coverage of low-resource languages.
@@
Mistral-7B Incorrect A. 1.00 1.00 1.00
Correct A. 1.00 1.00 1.00
Qwen1.5-7B Incorrect A. 0.89 0.99 0.94
Correct A. 0.95 0.67 0.78
Llama-2-7b Incorrect A. 0.90 1.00 0.95
Correct A. 1.00 0.70 0.83
Mixtral-8x7B Incorrect A. 0.85 0.98 0.91
Correct A. 0.97 0.76 0.85
Table 2: Evaluation of the GPT-4 classifier that we use to assess answer accuracy on MultiQ. Precision (P), Recall (R), and F1-Score of the GPT-4 judgments compared to human evaluation on answers from four different models to the same 100 random questions.
@@ 3.3 Question Answering Accuracy
To evaluate the reliability of our automated evaluation method, we tasked two independent human annotators to label model responses from four of our models for the same 100 randomly selected prompts as correct or incorrect. [^14] We exclude the 13B and 70B versions of Llama2 from the QA accuracy analysis for reasons of clarity, and because they resemble the results of Llama2 7B. Disagreements between the annotators, which occurred for no more than four responses per model, were resolved in discussions with one of the authors. We find that the accuracy of our automated evaluation, as measured against the human labels, is very high across models (see Table 2). Accuracy is perfect on Mistral. For the other models, the automated evaluation tends to be very precise on correct answers, but less so on incorrect answers. This means that the automated evaluation will likely underestimate the proportion of correct model answers across languages. Overall, automated evaluation, like automated translation, introduces noise to our silver standard MultiQ benchmark, but we find the amount of noise to likely be small.
@@ 3.4 Language Fidelity vs Answer Accuracy
Finally, we combine the results of §3.2 and §3.3 to assess the relationship between language fidelity and QA accuracy. Table 4 shows the mean accuracy grouped by response language category. We find that Llama2, despite its overall low language fidelity, shows strong QA accuracy. Especially when answering in the same language as prompted, it gives a correct answer in almost half of the cases (44.7%). Scaling the model further improves accuracy, with the 13B and 70B variants achieving higher correct answer rates of 51.6% and 61.4%, respectively, when answering in the same language as prompted. Qwen and Mistral show a similar pattern, i.e.  higher answering accuracy when answering in the prompt language with 21.5% and 17.1%. However, their overall accuracy is quite low, especially when compared to their initial high language fidelity. Strikingly, we find significant differences across models that ought to be similar due to their intended use for English, i.e.  Llama2 demonstrating low fidelity yet high accuracy, in contrast to Mistral, which exhibits the reverse pattern. A qualitative analysis of Mistral’s answers reveals that the model often merely repeats the questions it was asked, which is technically faithful but never an accurate answer. Only for Mixtral is this pattern not apparent, with the highest accuracy of 37% being achieved for responses in English. Overall, models answering in the prompted language tend to be more accurate than those responding in English, Mixtral being the exception. Therefore, our results suggest that increased language fidelity may positively impact QA accuracy.
4 Drivers of Multilinguality

### v2

@@ Abstract
Large language models (LLMs) should benefit everyone, including a global majority of non-English speakers. However, most LLMs today, and open LLMs in particular, are often intended for use in just English (e.g. Llama2, Mistral) or a small handful of high-resource languages (e.g. Mixtral, Qwen). Recent research shows that, despite limits in their intended use, people prompt LLMs in many different languages. Therefore, in this paper, we investigate the basic multilingual capabilities of state-of-the-art open LLMs beyond their intended use. For this purpose, we introduce MultiQ, a new silver standard benchmark for basic open-ended question answering with 27.4k test questions across a typologically diverse set of 137 languages. With MultiQ, we evaluate language fidelity, i.e. whether models respond in the prompted language, and question answering accuracy. All LLMs we test respond faithfully and/or accurately for at least some languages beyond their intended use. Most models are more accurate when they respond faithfully. However, differences across models are large, and there is a long tail of languages where models are neither accurate nor faithful. We explore differences in tokenization as a potential explanation for our findings, identifying possible correlations that warrant further investigation.
@@ 1 Introduction
Languages other than English remain underrepresented and underserved by state-of-the-art language technologies, posing a barrier to equal and inclusive AI (Bender, 2011; Joshi et al., 2020). While proprietary large language models (LLMs) like GPT-4 (OpenAI, 2023) may answer questions and follow instructions in many different languages, even the best and most popular open LLMs are much more restricted in their language coverage: Llama2-chat (Touvron et al., 2023), for example, is “intended for commercial and research use in English”. [^1] https://huggingface.co/meta-llama/Llama-2-7b-hf Yi is “bilingual” in English and Chinese [^2] https://huggingface.co/01-ai/Yi-34B-Chat, and Mistral-7b-instruct (Jiang et al., 2023) “only works in English”. [^3] https://mistral.ai/news/la-plateforme/
@@
Even though most open LLMs are restricted in their intended use to one or a handful of languages, datasets of real-world LLM usage show that people prompt LLMs in many different languages, often beyond their intended use (Ouyang et al., 2023; Zhao et al., 2024; Zheng et al., 2024). This has motivated initial research into the multilingual capabilities of monolingual models (Armengol-Estapé et al., 2022; Lai et al., 2023). However, this research has mostly focused on older proprietary LLMs and on a relatively small number of languages and/or specific tasks.
In this paper, we investigate the basic multilingual capabilities of a variety of state-of-the-art chat-optimized open LLMs across a typologically diverse set of 137 languages. Specifically, we ask two main research questions that correspond to two dimensions of multilingual capability: 1) What is the multilingual language fidelity of current chat-optimized open LLMs, and 2) What is the multilingual question answering (QA) accuracy of current chat-optimised open LLMs? Language fidelity describes the ability to respond to prompts in the prompted language. QA accuracy describes the ability to give correct answers to open-ended questions, in the prompted language or any other. An ideal multilingual model would give answers that are both faithful and correct.
@@
    Positive Interaction: Increased language fidelity appears to positively impact answer accuracy, since model answers that match the prompt language tend to be more accurate.
@@
    Tokenization as (Partial) Explanation: Models tend to achieve higher accuracy on languages they can tokenize into subwords instead of characters or ASCII tokens.
We publish all data and code at https://github.com/paul-rottger/multiq
@@ 2.1 Dataset Creation
For the initial English questions, we used two different sources to increase question diversity. a) We collected 100 questions from the LMSYS-Chat-1M dataset (Zheng et al., 2024), which catalogs real-world user interactions with LLMs. Specifically, we sampled all single-sentence English-language and sorted them by frequency. Then, we manually selected from the top until we reached 100 questions. This portion of our data directly reflects real-world LLM usage. b) We manually created another set of 100 questions evenly spread across 10 different subjects at elementary to middle school level (e.g. mathematics and geography). To maximize the diversity of the questions, we prompted GPT-4 to provide us with a set of simple and clear questions with simple and clear answers for each of the subjects. We then iterated and manually selected questions until we reached 10 questions per subject. This portion of our data expands MultiQ’s topical coverage.
@@
Source Example Prompts
LMSYS Was the year 2000 a leap year?
What is 2 + 2 * 3?
How many feet does a chicken have?
@@
Table 1: Examples of English questions covered in MultiQ. We present three prompts from each source (LMSYS and GPT-4) covering different domains.
@@ 2.2 Validation of Translations
We asked native speakers to annotate the correctness of the 200 translated MultiQ questions for 19 languages: Arabic, Catalan, Chinese, Farsi, French, German, Hindi, Indonesian, Italian, Japanese, Korean, Spanish, Tagalog, Russian, Spanish, Quechua, Ukrain, Urdu and Xhosa. Across these languages, annotators marked an average of 91.6% of translations as correct. Translations were least accurate for Tagalog at 60.0% and Arabic at 82.2%, while the translations for Italian and German were the most accurate, at 99.0%. We present the full results in the Appendix A. Qualitatively, several annotators stated that some translations, while accurate in content, tended to be literal, word-for-word translations rather than natural expressions. Overall, the automated translation introduces some noise into MultiQ, but our validation results suggest that the amount of noise is limited. This is why we frame MultiQ as a silver standard benchmark that can provide meaningful insights into basic multilingual capabilities, even if exact results on individual test cases may not be perfectly reliable.
@@ 2.3 Typological Diversity
To estimate language distances, we use the lang2vec toolkit (Littell et al., 2017) which contains precomputed language distances based on the typological language information queried from the URIEL knowledge base. [^5] https://www.cs.cmu.edu/~dmortens/projects/7_project/ Following Ploeger et al. (2024), we calculate the distances between all pairs of languages in our dataset that have at least 5% coverage in the URIEL vectors. Figure 2 shows the distribution of pairwise geographic, syntactic, and genetic distances of all covered languages. We find that the languages in MultiQ cover a wide range of typologically similar and distant language pairs with an expected high skewness of genetic distance complementing previous research (Ploeger et al., 2024). We can therefore confidently speak of a high typological diversity of our dataset.
@@
Next, we calculate the typological feature coverage of the 137 languages in MultiQ, using language features provided in the Grambank database (Skirgård et al., 2023). For this purpose, we map the language IDs in MultiQ (obtained from the Google Translate API) to the Glottoids used by Grambank. [^6] For 14 languages in our dataset there is no matching entry in Grambank. To avoid incorrect mapping, we do not assign them manually and exclude these languages from the calculation Taken together, the languages in MultiQ cover at least 95.4% of the typological features recoded in Grambank. This underlines the typological diversity of our dataset.
@@ 3 Experiments and Results
Using MultiQ, we can now answer our two main research questions regarding the multilingual language fidelity and multilingual QA accuracy of current chat-optimized open LLMs. We first separately assess fidelity and accuracy, and then evaluate their relationship.
@@ 3.1 Overall Experimental Setup
We test six open-access LLMs that are both popular and competitive in performance with other state-of-the-art models as measured on standard (English-language) benchmarks such as the LMSys Leaderboard [^9] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard and AlpacaEval. [^10] https://tatsu-lab.github.io/alpaca_eval/ Specifically, the 7B, 13B and 70B versions of Llama2-Chat (Touvron et al., 2023), the 7B Mistral-Instruct-v0.1 (Jiang et al., 2023), the 8x7B Mixtral-Instruct-v0.1 (Jiang et al., 2024) and the 7B Qwen1.5-Chat model (Bai et al., 2023). We test three sizes of Llama2 to evaluate scaling on MultiQ. Llama2 and Mistral are explicitly intended for English use only, whereas Mixtral and Qwen are explicitly multilingual: Mixtral “handles English, French, Italian, German and Spanish”, while Qwen offers unspecified “multilingual support”.
@@ 3.2 Language Fidelity
To confirm the robustness of our findings, we investigate the impact of brief and numerical responses from the models on our results. To this end, we excluded all questions from the MultiQ dataset that required a numerical answer, specifically removing the 10 curated questions of the domain “math” as well as 16 questions that were drawn from the LMSYS dataset. We then analyzed the character length of the model responses, noting an average length of 270-670 characters across different models. We also exclude responses shorter than 10 characters from the language fidelity calculation. The overall language fidelity of the models showed only marginal changes in response to these changes, underscoring the minimal influence of brief or numerical responses on our analysis.
@@
Next, we conduct a more fine-grained analysis of model answers in “another language”, i.e. neither the input language nor English. For this, we focus on the three models responding in another language more than 10% of the time (Qwen, Mistral, Mixtral). Of these, Mistral demonstrates the highest level of diversity among the languages it responds in. GlotLID identifies more than 360 distinct languages in Mistral’s responses, compared to around 240 languages for Mixtral and Qwen. We also examine the correspondence between the language family of input and output language. While Mixtral provides a response in the same language family at least 68.7% of the time, if not in an exact match or responding in English, Qwen demonstrates this behavior only 60.6% of the time, while Mistral does so for 54.6% of prompts. Note, however, that for 6% of the answers of Mistral and 7% of the answers of Mixtral, the language family of the response languages could not be determined and is thus classified as ’unknown’. Additionally, the distribution of language families in MultiQ is highly skewed, with 45.3% of the languages belonging to the ’Indo-European’ family, which spans a very broad range of languages from Irish to Turkish.
Finally, we analyze the relationship between the language of the prompt and the frequency of the respective response language of the models. Figure 4 shows correlation matrices for Mistral and Qwen. Corresponding matrices for the other models are shown in Appendix C. We find that Hindi is the most frequently selected language outside of the input language or English, closely followed by Indonesian. For Qwen, for example, these languages make up 20.4% of the “another language” category. This is visible by the thin yellow vertical lines leading to the axis ticks ind and hin in the Figure for both models respectively. A potential explanation may be that the models lack support for numerous languages from India and Indonesia, thus treating Hindi and Indonesian as some kind of ’fallback’ languages for the wider language area. For example, the models often respond to languages such as Malay and Javanese in Indonesian, and languages such as Maithili, Konkani, and Bhojpuri in Hindi. We also observe this phenomenon for smaller European languages. For example, the three models respond to questions in Croatian mostly in Italian, to those in Luxembourgish mostly in German, and those in Galician mostly in Portuguese. These observations underline the importance of improving multilingual models in the language coverage of low-resource languages.
@@
Mistral-7B Incorrect A. 0.97 1.00 0.98
Correct A. 0.98 0.87 0.92
Qwen1.5-7B Incorrect A. 0.88 0.99 0.93
Correct A. 0.94 0.61 0.74
Llama-2-7b Incorrect A. 0.92 1.00 0.96
Correct A. 1.00 0.74 0.85
Mixtral-8x7B Incorrect A. 0.82 0.98 0.89
Correct A. 0.97 0.71 0.82
Table 2: Evaluation of the GPT-4 classifier that we use to assess answer accuracy on MultiQ. We show Precision (P), Recall (R), and F1-Score of GPT-4 judgments on responses from four models to the same sample of 282 questions covering all languages, which were annotated by humans for whether they are correct or not.
@@ 3.3 Question Answering Accuracy
To evaluate the reliability of our automated evaluation method, we tasked two independent human annotators to label model responses from four of our models for the same 282 randomly selected prompts covering at least two questions per language, as correct or incorrect. [^13] We exclude the 13B and 70B versions of Llama2 from the QA accuracy analysis for reasons of clarity, and because they resemble the results of Llama2 7B. Disagreements between the annotators, which occurred for no more than four responses per model, were resolved in discussions with one of the authors. We find that the accuracy of our automated evaluation, as measured against the human labels, is very high across models (see Table 2). For all models, the automated evaluation tends to be very precise on correct answers, but less so on incorrect answers. This means that the automated evaluation will likely underestimate the proportion of correct model answers across languages. Overall, automated evaluation, like automated translation, introduces noise to our silver standard MultiQ benchmark, but we find the amount of noise to likely be small.
@@ 3.4 Language Fidelity vs Answer Accuracy
Finally, we combine the results of §3.2 and §3.3 to assess the relationship between language fidelity and QA accuracy. Table 4 shows the mean accuracy grouped by response language category. We find that Llama2, despite its overall low language fidelity, shows strong QA accuracy. Especially when answering in the same language as prompted, it gives a correct answer in almost half of the cases (44.7%). Scaling the model further improves accuracy, with the 13B and 70B variants achieving higher correct answer rates of 51.6% and 61.4%, respectively, when answering in the same language as prompted. Qwen and Mistral show a similar pattern, i.e. higher answering accuracy when answering in the prompt language with 21.5% and 17.1%. However, their overall accuracy is quite low, especially when compared to their initial high language fidelity. Strikingly, we find significant differences across models that ought to be similar due to their intended use for English, i.e. Llama2 demonstrating low fidelity yet high accuracy, in contrast to Mistral, which exhibits the reverse pattern. A qualitative analysis of Mistral’s answers reveals that the model often merely repeats the questions it was asked, which is technically faithful but never an accurate answer (see Table 11). Only for Mixtral is this pattern not apparent, with the highest accuracy of 37% being achieved for responses in English. Overall, we find that if models answer in the same language as the prompted language, they tend to be more accurate than if they respond in English. The only exception of this pattern is Mixtral. Therefore, our results suggest that increased language fidelity may positively impact QA accuracy.
4 Tokenization and Multilinguality